OpenStudy (anonymous):

The cost of manufacturing x items is given by C(x) = 0.02x^2 + 20x + 1800. Find the marginal cost function. Compuare the marginal cost at x = 20 to the actual cost of producing the 20th item.

OpenStudy (anonymous):

This is what i did but my teacher says it wrong. C(x) = 0.02x^2 + 20x + 1800 C’(x) = 0.04x + 20 C’(20) = 0.04(20) + 20 = 20.8 so this will be the marginal cost at x = 20 Acutal cost of the 20th item : c(20) - c(19) = 0.02(20^2) + 20(20) + 1800 – 0.02(19)^2 + 20(19) + 1800 = 4380.78

OpenStudy (anonymous):

@agent0smith can u help?

OpenStudy (anonymous):

c(20)-c(19) = 2208 - 2187,22 = 20.78 it is just one algebraic mistake

OpenStudy (anonymous):

okay thanks so thats the marginal cost?

OpenStudy (anonymous):

no! u already got the amrginal cost "C’(20) = 0.04(20) + 20 = 20.8 so this will be the marginal cost at x = 20 "

OpenStudy (anonymous):

ohh okay so what do i do next?

OpenStudy (anonymous):

u wrote this "Acutal cost of the 20th item : c(20) - c(19) = 0.02(20^2) + 20(20) + 1800 – 0.02(19)^2 + 20(19) + 1800 = 4380.78 " i m telling u that part is wrong with algebraic mistakes

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so you're saying in that its where i made my sitake?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i made that calculations and i got c(20)-c(19) = 2208 - 2187,22 = 20.78

OpenStudy (anonymous):

so how about the actual cost?

OpenStudy (anonymous):

@Loser66 so how would i find the actual cost? should i just substitute 20 into the first problem?

OpenStudy (loser66):

sorry friend, I totally lost you in this question. I don't get why you subtract C(20 ) -C(19)

OpenStudy (anonymous):

to get the marginal cost? i'm so confused :/

OpenStudy (loser66):

@ganeshie8

OpenStudy (loser66):

@RadEn

OpenStudy (loser66):

@mathstudent55

OpenStudy (loser66):

I know why don't get you, hihi, it's economic problem. hand off, friend. never known before

OpenStudy (anonymous):

hi again so what part u dont understand i think the problem is done

OpenStudy (anonymous):

i'm trying to figure it out what is the actual cost is? since we already found marginal cost

OpenStudy (anonymous):

u have the marginal cost C’(x) = 0.04x + 20 C’(20) = 0.04(20) + 20 = 20.8 and the actual cost of producing the 20th item. c(20)-c(19) = 2208 - 2187,22 = 20.78 the problem says that u must compare them and u can can see they r very close

OpenStudy (anonymous):

how how do i do that?

OpenStudy (anonymous):

how to do what? compare them? 20.8 is very close to 20.78 u dont need to do anymore

OpenStudy (anonymous):

okay thanks